OpenStudy (anonymous):

Given f(x) = 3x + 7, and g(x) = x − 1 ƒ(g(5)) =3x+32 right?

jimthompson5910 (jim_thompson5910):

f(g(5)) should be a constant number

jimthompson5910 (jim_thompson5910):

ie there shouldn't be any variables in the answer

jimthompson5910 (jim_thompson5910):

what is the value of g(5)

OpenStudy (anonymous):

19 dear :)

jimthompson5910 (jim_thompson5910):

pengembara_bumi1 you're not helping anyone by just giving out answers

OpenStudy (anonymous):

is it 62?

jimthompson5910 (jim_thompson5910):

g(x) = x − 1 g(5) = 5 − 1 ... replace every x with 5 g(5) = ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

now that we know that g(5) = 4, we can use this to find f(g(5))

OpenStudy (anonymous):

22 for f?

jimthompson5910 (jim_thompson5910):

f(x) = 3x + 7 f(g(x)) = 3( g(x) ) + 7 ... replace every x with g(x) f(g(5)) = 3( g(5) ) + 7 ... replace every x with 5 f(g(5)) = 3( 4 ) + 7 ... replace g(5) with 4 (since we just found that g(5) = 4) f(g(5)) = 12 + 7 f(g(5)) = 19

OpenStudy (anonymous):

fg(x)=? when x=5 g(5)=5-1 g(5)=4 f(4)=3(4)+7 fg(5)=19

jimthompson5910 (jim_thompson5910):

hopefully all that makes sense

OpenStudy (anonymous):

sorry @jim_thompson5910 :(

jimthompson5910 (jim_thompson5910):

its fine, just keep it in mind

OpenStudy (anonymous):

thank you!!! both of you!!!!

jimthompson5910 (jim_thompson5910):

you're welcome